For the following set of cash flows,

  

Year Cash Flow
0 –$7,500           
1 3,500           
2 6,200           
3.

3,700

d. What is the NPV at a discount rate of 24 percent?
Answer #1
Year Cash flow
0 -7500
1 3500
2 6200
3 3700

The cash flows are:

C0 = -7500, C1 = 3500, C2 = 6200, C3 = 3700

Discount rate = r = 24%

NPV is calculated using the formula:

NPV = C0 + C1/(1+r)1 + C2/(1+r)2 + C3/(1+r)3

NPV = -7500 + 3500/(1+24%)1 + 6200/(1+24%)2 + 3700/(1+24%)3 = -7500 + 2822.58064516129 + 4032.25806451613 + 1940.60286663758 = 1295.44157631499

NPV = $1295.44 (Rounded to two decimals)

Answer -> NPV = $1295.44